# Evolution of a Single Impurity Across the Superfluid–Mott insulator Transition in the Bose–Hubbard Model

### v1

@@ Abstract
We investigate how the coherence and spatial dressing of a single impurity evolve in the two-dimensional Bose–Hubbard model when the impurity couples attractively to the bath. Using large-scale, sign-problem-free quantum Monte Carlo simulations based on the worm algorithm, we track the impurity winding number, the bath superfluid density and compressibility, and impurity–bath density correlations. First, by fixing the bath interaction at $U_{\mathrm{b}}/t=13.3$ in the superfluid regime and tuning the attractive impurity-bath coupling from $U_{\mathrm{ib}}/t=-1.0$ to 40.0-40.0, we uncover an interaction-driven winding-collapse localization: the impurity evolves from a mobile light polaron with finite winding to a heavy polaron and, finally, to a bound cluster with vanishing winding, while the bath remains globally superfluid. Second, we analyze impurity in Mott-insulating baths for both attractive and repulsive impurity–bath couplings, contrasting the resulting deformation clouds and localization patterns. Third, for a moderate attractive impurity-bath coupling $U_{\mathrm{ib}}/t=-8.0$, we tune the bath interaction $U_{\mathrm{b}}/t$ across the superfluid–Mott-insulator transition. The impurity evolves from a mobile light polaron with an extended deformation cloud in the superfluid to a nearly free defect propagating in an incompressible Mott-insulating background; for this moderate attraction it does not develop into a fully localized defect with a strictly local dressing cloud, even deep in the Mott insulator phase. This evolution constitutes a compressibility-controlled localization crossover of coherent impurity motion. Finally, together with our companion Letter [The Fate of a Single Impurity in the Bose–Hubbard Model], which focused on repulsive impurity-bath couplings, these results provide a unified microscopic picture of impurity localization in the Bose–Hubbard model, connecting interaction-driven and compressibility-controlled mechanisms across both attractive and repulsive regimes.
@@ I Introduction
The behavior of a single impurity immersed in a quantum many-body medium provides a sensitive probe of correlations, coherence, and emergent quasiparticles. In weakly interacting Bose gases, such an impurity forms a Bose polaron: a dressed quasiparticle characterized by a finite quasiparticle residue, renormalized effective mass, and an extended bath density cloud [1, 2, 3, 3, 4, 5, 6, 7, 8, 9]. In the continuum, Bose polarons have been observed and characterized with high precision using radio-frequency and Ramsey spectroscopy in ultracold atomic mixtures [10, 11, 12], and their properties across weak to strong coupling have been mapped out both experimentally and theoretically [1]. These studies have established the continuum Bose polaron as a paradigmatic example of an impurity problem in a quantum fluid.
@@
Beyond continuum Bose gases, optical lattices offer a complementary platform where the bath is naturally described by the Bose–Hubbard model (BHM) [13, 14, 15, 16], providing a controlled setting to study a single mobile impurity in a strongly correlated lattice Bose fluid. By tuning the ratio between on-site repulsion and tunneling, the BHM realizes a quantum phase transition from a compressible superfluid (SF) to an incompressible Mott insulator (MI), as first observed in pioneering cold-atom experiments [17]. With the advent of quantum gas microscopy, site-resolved imaging and control of individual atoms in optical lattices has become possible [18, 19], including the preparation and tracking of single impurity or spin excitation in Bose–Hubbard chains and 2D lattices [20]. This opens the door to directly visualizing how an impurity propagates, dresses, and eventually localizes as the background evolves from SF to MI.
@@
Theoretical work on lattice Bose polarons and impurities in the BHM has addressed a number of important limits [21, 22, 23, 24, 25, 26]. Variational approaches have explored mobile impurities near the SF–MI transition [24], and recent large-scale quantum Monte Carlo (QMC) simulations have analyzed the “Bose–Hubbard polaron” from weak to strong impurity–bath coupling in two dimensions [27]. In particular, Ref. [27] used worm-algorithm QMC in the grand-canonical ensemble to show that, at weak coupling, polaron spectral properties such as the effective mass can serve as a sensitive probe of the MI–SF transition, while at strong coupling the impurity binds an extra particle or hole in the MI, leading to novel bound states beyond the simple polaron picture. Nevertheless, most lattice studies to date have focused primarily on impurity spectral observables—energy shifts, effective masses, and spectral functions—and typically emphasized either the SF side or the MI side separately, with the bath entering only through averaged properties. What is still lacking is a unified, real-space characterization that follows both the impurity’s coherent mobility and the associated impurity-centered bath density deformations across the superfluid–Mott insulator transition.
@@
In a recent companion Letter [28], we addressed the fate of a single mobile impurity in the two-dimensional Bose–Hubbard model at unit filling by mapping out its regimes in the plane of bath interaction $U_{\mathrm{b}}/t$ and local impurity–bath coupling $U_{\mathrm{ib}}/t$ [28]. Using sign-problem-free multi-species worm-algorithm quantum Monte Carlo (QMC) simulations in the grand-canonical ensemble [29, 30, Lingua_2018], we demonstrated that impurity localization in this setting can be realized via two distinct yet unified routes. (i) Along lines of varying impurity-bath coupling $U_{\mathrm{ib}}/t$ at fixed bath parameters in a SF bath, localization is driven by an interaction-driven winding-collapse crossover: a weakly coupled impurity forms a mobile light polaron with finite winding and an extended dressing cloud, whereas increasing $\lvert U_{\mathrm{ib}}\rvert/t$ leads to the collapse of impurity winding while the bath remains globally superfluid, signaling a heavy self-trapped state and, for repulsive $U_{\mathrm{ib}}/t>0$, finally a depletion saturated-bubble state. (ii) Along lines of varying bath interaction $U_{\mathrm{b}}/t$, where the bath evolves from a SF to a MI state at fixed impurity-bath coupling $U_{\mathrm{ib}}/t$, the impurity localization is instead compressibility-controlled. Starting from a compressible SF bath, the impurity appears as a mobile light polaron with finite winding and an extended deformation cloud. As $U_{\mathrm{b}}/t$ increases, the bath superfluid density and compressibility are progressively suppressed and the impurity winding is continuously reduced. Upon entering the MI regime, the impurity crosses over to an almost free defect moving in a weakly distorted incompressible background with nearly flat deformation cloud, and in the deep MI phase its deformation cloud shrinks to within roughly two or three nearby sites, yielding a localized vacancy- or particle-like defect. This evolution is reflected in the impurity-centered cumulative bath density deformation $\Delta N(R)$, which changes from a broad polaronic profile in the SF to an almost flat response upon entering the MI, and eventually to a defect in the deep MI. In this way, the Letter established a global $U_{\mathrm{b}}/t$$U_{\mathrm{ib}}/t$ phase diagram that connects light polarons, heavy polarons, and saturated-bubble/bound-cluster regimes for both signs of $U_{\mathrm{ib}}/t$ in the SF bath, as well as nearly free and fully localized defect states in the MI bath. It further clarified that, along interaction-driven trajectories in the SF, impurity localization occurs via a continuous crossover rather than a thermodynamic phase transition, whereas in the MI region it can become a sharp transition from a nearly free particle to a localized defect state, while at fixed $U_{\mathrm{ib}}/t$ the evolution of the impurity across the SF–MI boundary remains a smooth, compressibility-controlled crossover.
@@
While the Letter [28] establishes a broad global picture of impurity localization in the Bose–Hubbard model, it does not fully develop a microscopic description of impurity behavior for attractive impurity–bath couplings, nor of impurity dynamics in genuinely incompressible Mott-insulating (MI) baths. In the MI regime of the Bose–Hubbard model, long-wavelength density fluctuations are gapped and the superfluid stiffness vanishes [CapogrossoSansonePRA2008]. At weak $U_{\mathrm{ib}}/t$, an impurity is expected to propagate almost as a free defect on top of a rigid background, while at strong $\lvert U_{\mathrm{ib}}\rvert/t$ it should nucleate a localized particle- or hole-like defect, as suggested by Ref. [27]. The global phase diagram in our letter already indicated where interaction-driven winding-collapse and compressibility-controlled localization lines meet inside the MI region, but several key microscopic questions were left open: (i) How does the interaction-driven winding-collapse mechanism operate microscopically for attractive impurity–bath couplings in a superfluid background? (ii) Once the bath is strictly incompressible, to what extent do attractive and repulsive impurities exhibit qualitatively different localization patterns and deformation clouds? (iii) For a fixed attractive impurity-bath coupling, how does the compressibility-controlled localization proceed as the bath is tuned from the superfluid into the Mott-insulating regime?
@@
In this paper, we build directly on the letter [28] and address these questions by performing a systematic, numerically exact QMC study that focuses on attractive impurity–bath couplings and provides a detailed account of impurity behavior in Mott-insulating baths. We analyze three complementary trajectories. (i) In a compressible SF bath, we implement an interaction-driven winding-collapse protocol for attractive impurities by fixing the bath interaction at $U_{\mathrm{b}}/t=13.33$ and tuning $U_{\mathrm{ib}}/t$ from 1.0-1.0 to 40.0-40.0, thereby following the evolution from a mobile light polaron with finite winding to a heavy polaron and then to a bound cluster while the bath remains globally superfluid. (ii) Inside the MI regime, we systematically compare attractive and repulsive impurity–bath couplings, revealing how nearly free defects and localized particle- or vacancy-like defects emerge as distinct localization patterns in an incompressible background. (iii) To expose the role of bath compressibility for attractive impurity, we implement a compressibility-driven localization protocol by fixing $U_{\mathrm{ib}}/t=-8.0$ and tuning $U_{\mathrm{b}}/t$ from the deep SF regime, through the critical region, into the deep MI, tracking how the impurity winding and deformation cloud collapse as the bath turns incompressible. For each parameter set we measure the impurity winding number, the bath superfluid response and compressibility, as well as radial impurity-centered real-space correlator $C(R)$ and the cumulative bath-density deformation $\Delta N(R)$. Taken together with our companion Letter [28], these results turn the global picture of interaction-driven and compressibility-controlled localization into a fully microscopic description that covers both superfluid and Mott-insulating baths and both attractive and repulsive impurity–bath couplings.
@@
Our results complement and extend previous work in several ways. First, by combining worldline winding diagnostics with radical impurity-centered real-space correlator $C(R)$ and cumulative bath-density deformation $\Delta N(R)$, we provide a direct visualization of the impurity’s localization process in both compressible SF and incompressible MI backgrounds, rather than inferring it solely from effective masses or ground state energy. Second, we clarify how interaction-driven winding collapse and compressibility-controlled localization operate at the microscopic level for attractive impurities: the same winding-based criterion, together with the deformation cloud, tracks the evolution from a mobile light polaron to a heavy polaron and finally to a bound cluster in the SF, and its subsequent modification as the bath loses compressibility. Third, by following the impurity across the SF–MI transition and into the deep MI regime, we show how the light polaron identified in the SF continuously evolves into a nearly free defect and then into a localized particle- or vacancy-like defect, thereby tying together the polaronic and defect descriptions within a single microscopic framework. Given the rapid experimental progress in site-resolved control of impurities in optical lattices [18, 19, 20], we expect that the observables and regimes identified here can be probed directly in future quantum-gas-microscope experiments.
@@
The remainder of this paper is organized as follows. Section II introduces the two-component Bose–Hubbard Hamiltonian, and Sec. III describes the quantum Monte Carlo method and the definitions of single-impurity properties, bath observables, and impurity-centered quantities such as the impurity-centered correlation $C_{\mathrm{ib}}(\mathbf{r})$ and the cumulative bath-density deformation $\Delta N(R)$. Sections IV, V, and VI present our main numerical results. We first briefly summarize the overall impurity phase diagram previously established in [28], and then analyze in detail three complementary trajectories: (i) an interaction-driven winding-collapse path in a SF bath with attractive impurity-bath coupling $U_{\mathrm{ib}}/t$, (ii) impurity physics in a Mott-insulating bath for representative attractive and repulsive couplings $U_{\mathrm{ib}}/t=\pm 8.0$, and (iii) a compressibility-driven localization path where the bath is tuned from SF to MI at fixed $U_{\mathrm{ib}}/t=-8.0$. For each case, we compare the evolution of the impurity winding number, the bath superfluid density response, and the impurity-centered real-space correlator $C(R)$ and its cumulative bath-density deformation $\Delta N(R)$. A summary of the main conclusions and an outlook are given in Sec. IX.
@@ II Model
where $b_{i}^{\dagger}$ ($a_{i}^{\dagger}$) creates a bath (impurity) boson on lattice site $i$, and $n_{\mathrm{b},i}=b_{i}^{\dagger}b_{i}$, $n_{\mathrm{imp},i}=a_{i}^{\dagger}a_{i}$ are the corresponding on-site number operators. The bath hopping amplitude $t_{\mathrm{b}}$ sets the energy scale ($t_{\mathrm{b}}=1$ throughout), and the on-site repulsion $U_{\mathrm{b}}$ controls the superfluid–Mott insulator transition of the bath. The impurity hopping $t_{\mathrm{imp}}=t=1$ is taken equal to $t_{\mathrm{b}}$. The local impurity–bath coupling $U_{\mathrm{ib}}$ is the key tuning parameter: $U_{\mathrm{ib}}>0$ corresponds to repulsion, $U_{\mathrm{ib}}<0$ to attraction. The bath chemical potential $\mu_{\mathrm{b}}$ is adjusted to realize unit filling, while a single impurity is enforced by $\sum_{i}\langle n_{\mathrm{imp},i}\rangle=1$.
Physical regimes.— Figure 1 illustrates the contrast of a single impurity between a compressible superfluid (SF) bath and an incompressible Mott-insulating (MI) bath. In the SF regime, bath bosons (blue wavy lines) form a coherent condensate with long-range phase conhenrence. A repulsive impurity (red disk) pushes away nearby bath particles, creating a broad depletion cloud, while an attractive impurity draws in additional bath density, forming a smooth accumulation halo. Because the bath is compressible, these density modulations extend over many sites and the impurity exhibits coherent motion with finite winding. In the MI regime, density fluctuations are frozen and only local distortions survive. At smaller $U_{\mathrm{b}}/t$, the impurity acts as a nearly-free defect for both repulsive and attractive impurity-bath couplings $U_{\mathrm{ib}}/t$, while at larger $U_{\mathrm{b}}/t$ it turns into a localized defect: a vacancy defect on the repulsive side and a particle defect on the attractive side, both characterized by vanishing impurity winding.
@@
Figure 1: Schematic impurity behavior in superfluid (SF) and Mott-insulating (MI) baths. The bath undergoes a SF–MI transition at $U_{\mathrm{b}}/t\simeq 16.7$. (a) In a compressible SF bath, bosons (blue wavy background) possess long-range phase coherence and support extended density fluctuations. A repulsive impurity (red disk) creates a broad depletion cloud, whereas an attractive impurity induces a smooth accumulation halo. Because the bath is compressible, these density modulations spread over many sites and the impurity retains coherent motion with finite winding, forming a light polaron at weak coupling and an extended heavy state at stronger coupling. (b) In an incompressible MI bath, density fluctuations are frozen and only short-range particle–hole excitations are allowed. For both repulsive and attractive impurity–bath couplings, the impurity behaves as a nearly free defect at moderate $U_{\mathrm{b}}/t$, producing only weak local distortion. At larger $U_{\mathrm{b}}/t$ the impurity becomes fully localized: a repulsive impurity expels one bath particle and forms a quantized vacancy defect, while an attractive impurity binds extra one boson into a particle defect. Both MI-side defects exhibit short-range density distortion and vanishing impurity winding, in stark contrast to the extended polaronic dressing in the SF regime.

### v2

# Mobile impurity coupled to correlated lattice bosons

@@ Abstract
We investigate how the coherence and spatial dressing of a single impurity evolve in the two-dimensional Bose–Hubbard model when the impurity couples attractively to the bath. Using large-scale, sign-problem-free worm-algorithm quantum Monte Carlo, we measure the impurity winding, bath superfluid response and compressibility, and impurity–bath density correlations. In a compressible superfluid bath ($U_{\mathrm{b}}/t=13.3$), strengthening attraction drives an interaction-controlled winding-collapse self-trapping crossover: a mobile light polaron evolves continuously into a heavy polaron and ultimately into a bound cluster with vanishing winding, while the bath remains globally superfluid. In incompressible Mott-insulating baths, by contrast, extended density rearrangements are suppressed and the dressing cloud collapses; we compare the resulting short-range deformation patterns for both attractive and repulsive couplings. Across the SF–MI transition at fixed moderate attraction ($U_{\mathrm{ib}}/t=-8.0$), the impurity crosses over from a mobile polaron with an extended deformation cloud to a nearly free defect with minimal dressing in the Mott background; for this coupling it does not lock into a fully self-trapped defect even deep in the insulator. Together with our companion Letter [Impurity Self-Trapping in Lattice Bose systems] on repulsive couplings, these results provide a unified microscopic picture of impurity self-trapping in correlated lattice bosons, connecting interaction-driven winding collapse in the superfluid to compressibility-controlled undressing across the SF–MI transition.
@@ I Introduction
The behavior of a single impurity immersed in a quantum many-body medium provides a sensitive probe of correlations, coherence, and emergent quasiparticles. In weakly interacting Bose gases, such an impurity forms a Bose polaron: a dressed quasiparticle characterized by a finite quasiparticle residue, a renormalized effective mass, and an extended bath-density deformation cloud [1, 2, 3, 4, 5, 6, 7, 8, 9]. In the continuum, Bose polarons have been observed and characterized with high precision using radio-frequency and Ramsey spectroscopy in ultracold atomic mixtures [10, 11, 12], and their properties across weak to strong coupling have been mapped out both experimentally and theoretically [1]. These studies have established the continuum Bose polaron as a paradigmatic impurity problem in a quantum fluid.
@@
Beyond continuum Bose gases, optical lattices offer a complementary platform where the bath is naturally described by the Bose–Hubbard model (BHM) [13, 14, 15, 16], providing a controlled setting to study a single mobile impurity in a strongly correlated lattice Bose fluid. By tuning the ratio between on-site repulsion and tunneling, the BHM realizes a quantum phase transition from a compressible superfluid (SF) to an incompressible Mott insulator (MI), as first observed in pioneering cold-atom experiments [17]. With the advent of quantum gas microscopy, site-resolved imaging and control of individual atoms in optical lattices has become possible [18, 19], including the preparation and tracking of a single impurity or spin excitation in Bose–Hubbard chains and two-dimensional lattices [20]. This opens the door to directly visualizing how an impurity propagates, dresses, and eventually self-traps as the background evolves from SF to MI.
@@
Theoretical work on lattice Bose polarons and impurities in the BHM has addressed a number of important limits [21, 22, 23, 24, 25, 26]. Variational approaches have explored mobile impurities near the SF–MI transition [24], and recent large-scale quantum Monte Carlo (QMC) simulations have analyzed the “Bose polaron” from weak to strong impurity–bath coupling in two dimensions [27]. In particular, Ref. [27] used worm-algorithm QMC in the grand-canonical ensemble to show that, at weak coupling, polaron spectral properties such as the effective mass can serve as a sensitive probe of the MI–SF transition, while at strong coupling the impurity binds an extra particle or hole in the MI, leading to bound states beyond the simple polaron picture. Nevertheless, most lattice studies to date have focused primarily on impurity spectral observables—energy shifts, effective masses, and spectral functions—and typically emphasized either the SF side or the MI side separately, with the bath entering mainly through averaged properties. What is still lacking is a unified, real-space characterization that follows both the impurity’s coherent mobility and the associated impurity-centered bath-density deformations across the SF–MI transition.
@@
In a recent companion Letter [28], we addressed the fate of a single mobile impurity in the two-dimensional Bose–Hubbard model at unit filling by mapping out its regimes in the plane of bath interaction $U_{\mathrm{b}}/t$ and local impurity–bath coupling $U_{\mathrm{ib}}/t$. Using sign-problem-free multi-species worm-algorithm QMC simulations in the grand-canonical ensemble [29, 30, 31], we showed that impurity self-trapping in this setting is organized by two distinct routes. First, within a compressible SF bath, increasing $|U_{\mathrm{ib}}|/t$ drives an interaction-driven winding-collapse crossover: a weakly coupled impurity forms a mobile light polaron with finite winding and an extended dressing cloud, whereas stronger coupling suppresses winding while the bath remains globally superfluid, producing a heavy, self-trapped state and, for repulsive $U_{\mathrm{ib}}/t>0$, eventually a depletion saturated-bubble configuration. Second, along trajectories where the bath is tuned across the SF–MI transition at fixed $U_{\mathrm{ib}}/t$, the evolution becomes compressibility controlled. Starting from a compressible SF, the impurity appears as a mobile light polaron with an extended deformation cloud, while increasing $U_{\mathrm{b}}/t$ progressively suppresses the bath superfluid response and compressibility and reduces the impurity winding. Upon entering the MI regime, the polaronic dressing collapses and the impurity becomes nearly free in an incompressible background; deeper in the MI, strong coupling can lock the impurity to a vacancy- or particle-type defect with a strongly localized deformation cloud. These trends are captured directly by the impurity-centered cumulative bath-density deformation $\Delta N(R)$, which evolves from a broad polaronic profile in the SF to a nearly flat response in the MI and, at strong coupling, to a compact defect signature. In this way, the Letter established a global $U_{\mathrm{b}}/t$$U_{\mathrm{ib}}/t$ diagram connecting light polarons, heavy polarons, and saturated-bubble/bound-cluster regimes in the SF, as well as nearly free and defect-bound self-trapped states in the MI. It further clarified that self-trapping within the SF proceeds via a continuous crossover (without a bath phase transition), whereas in the MI the change between nearly free and defect-bound sectors can become sharp, while the evolution across the SF–MI boundary at fixed $U_{\mathrm{ib}}/t$ remains a smooth, compressibility-controlled crossover.
@@
While the Letter [28] establishes a broad global picture of impurity self-trapping in the Bose–Hubbard model, it does not fully develop a microscopic description for attractive impurity–bath couplings, nor for impurity behavior deep in a genuinely incompressible Mott-insulating (MI) bath. In the MI regime, long-wavelength density fluctuations are gapped and the superfluid stiffness vanishes [16]. At weak $U_{\mathrm{ib}}/t$, an impurity is therefore expected to behave nearly as a free defect on top of a rigid background, whereas at strong $|U_{\mathrm{ib}}|/t$ it can bind and pin a particle- or hole-like defect, as suggested in Ref. [27]. Although the global diagram in Ref. [28] already indicated where the interaction-driven winding-collapse route in the SF and the compressibility-controlled route across the SF–MI boundary meet inside the MI region, several key microscopic questions remained open: (i) How does the interaction-driven winding-collapse mechanism manifest for attractive impurity–bath couplings in a superfluid background? (ii) Once the bath is strictly incompressible, in what ways do attractive and repulsive impurities differ in their deformation clouds and defect-binding patterns? (iii) For a fixed attractive coupling, how does the compressibility-controlled self-trapping route proceed as the bath is tuned from SF into the MI?
@@
In this paper, we build on Ref. [28] and address these questions with a systematic, numerically exact worm-algorithm QMC study that focuses on attractive impurity–bath couplings and provides a detailed account of impurity behavior in MI baths. We analyze three complementary trajectories. (i) In a compressible SF bath, we implement an interaction-driven winding-collapse protocol for attractive impurities by fixing $U_{\mathrm{b}}/t=13.33$ and tuning $U_{\mathrm{ib}}/t$ from 1.0-1.0 to 40.0-40.0, thereby following the evolution from a mobile light polaron with finite winding to a heavy polaron and ultimately to a bound cluster while the bath remains globally superfluid. (ii) Inside the MI regime, we compare attractive and repulsive couplings and show how the impurity crosses over from a nearly free defect to a defect-bound, self-trapped state (particle- or vacancy-like) as $|U_{\mathrm{ib}}|/t$ increases in an incompressible background. (iii) To isolate the role of bath compressibility for an attractive impurity, we fix $U_{\mathrm{ib}}/t=-8.0$ and tune $U_{\mathrm{b}}/t$ from deep SF, through the critical region, into deep MI, tracking how impurity winding and the deformation cloud collapse as the bath becomes incompressible. For each parameter set we measure the impurity winding number, the bath superfluid response and compressibility, and impurity-centered real-space observables including the radial correlator $C_{\rm ib}(R)$ and the cumulative bath-density deformation $\Delta N(R)$. Together with Ref. [28], these results promote the global two-route picture into a fully microscopic account that spans both attractive and repulsive couplings and both compressible and incompressible bath regimes.
@@
Our results complement and extend previous work in several ways. First, by combining worldline winding diagnostics with impurity-centered real-space observables ($C_{\rm ib}(R)$ and $\Delta N(R)$), we directly visualize the self-trapping process in both SF and MI backgrounds, rather than inferring it solely from effective masses or ground-state energies. Second, we clarify how interaction-driven winding collapse and compressibility-controlled undressing operate microscopically for attractive impurities: the same winding-based criterion, together with the deformation cloud, tracks the evolution from a mobile light polaron to a heavy polaron and finally to a bound cluster in the SF, and its subsequent modification as the bath loses compressibility. Third, by following the impurity across the SF–MI transition and into the deep MI regime, we show how the light polaron in the SF evolves continuously into a nearly free defect and, at stronger coupling in the MI, into a defect-bound self-trapped state, thereby connecting the polaronic and defect descriptions within a single framework. Given the rapid experimental progress in site-resolved control of impurities in optical lattices [18, 19, 20], we expect that the observables and regimes identified here can be probed directly in future quantum-gas-microscope experiments.
@@
The remainder of this paper is organized as follows. Section II introduces the two-component Bose–Hubbard Hamiltonian, and Sec. III describes the QMC method and defines the impurity and bath observables, including the impurity-centered correlator $C_{\mathrm{ib}}(\mathbf{r})$ and the cumulative bath-density deformation $\Delta N(R)$. Sections IV, V, and VI present our main numerical results. We first briefly summarize the global impurity diagram established in Ref. [28], and then analyze in detail three trajectories: (i) an interaction-driven winding-collapse path in an SF bath with attractive $U_{\mathrm{ib}}/t$, (ii) impurity physics in an MI bath for representative attractive and repulsive couplings $U_{\mathrm{ib}}/t=\pm 8.0$, and (iii) a compressibility-controlled path where the bath is tuned from SF to MI at fixed $U_{\mathrm{ib}}/t=-8.0$. A summary of the main conclusions and an outlook are given in Sec. IX.
@@ II Model
where $b_{i}^{\dagger}$ ($a_{i}^{\dagger}$) creates a bath (impurity) boson on lattice site $i$, and $n_{\mathrm{b},i}=b_{i}^{\dagger}b_{i}$, $n_{\mathrm{imp},i}=a_{i}^{\dagger}a_{i}$ are the corresponding on-site number operators. The bath hopping amplitude $t_{\mathrm{b}}$ sets the energy scale ($t_{\mathrm{b}}=1$ throughout), and the on-site repulsion $U_{\mathrm{b}}$ controls the superfluid–Mott insulator transition of the bath. The impurity hopping $t_{\mathrm{imp}}=t=1$ is taken equal to $t_{\mathrm{b}}$. The local impurity–bath coupling $U_{\mathrm{ib}}$ is the key tuning parameter: $U_{\mathrm{ib}}>0$ corresponds to repulsion, $U_{\mathrm{ib}}<0$ to attraction. The bath chemical potential $\mu_{\mathrm{b}}$ is adjusted to realize unit filling of the bath, while a single impurity is enforced by $\sum_{i}\langle n_{\mathrm{imp},i}\rangle=1$.
Physical regimes.— Figure 1 illustrates the contrast of a single impurity between a compressible superfluid (SF) bath and an incompressible Mott-insulating (MI) bath. In the SF regime, bath bosons (blue wavy lines) form a coherent condensate with long-range phase coherence. A repulsive impurity (red disk) pushes away nearby bath particles, creating a broad depletion cloud, while an attractive impurity draws in additional bath density, forming a smooth accumulation halo. Because the bath is compressible, these density modulations extend over many sites and the impurity exhibits coherent motion with finite winding. In the MI regime, density fluctuations are frozen and only local distortions survive. Within the MI regime, for moderate $U_{\mathrm{b}}/t$ the impurity acts as a nearly-free defect for both repulsive and attractive impurity-bath couplings $U_{\mathrm{ib}}/t$, while deeper in the MI (larger $U_{\mathrm{b}}/t$) it turns into a self-trapped defect: a vacancy defect on the repulsive side and a particle defect on the attractive side, both characterized by vanishing impurity winding.
@@
Figure 1: Schematic impurity behavior in superfluid (SF) and Mott-insulating (MI) baths. The bath undergoes a SF–MI transition at $U_{\mathrm{b}}/t\simeq 16.7$. (a) In a compressible SF bath, bosons (blue wavy background) possess long-range phase coherence and support extended density fluctuations. A repulsive impurity (red disk) creates a broad depletion cloud, whereas an attractive impurity induces a smooth accumulation halo. Because the bath is compressible, these density modulations spread over many sites and the impurity retains coherent motion with finite winding, forming a light polaron at weak coupling and an extended heavy state at stronger coupling. (b) In an incompressible MI bath, density fluctuations are frozen and only short-range particle–hole excitations are allowed. For both repulsive and attractive impurity–bath couplings, the impurity behaves as a nearly free defect at moderate $U_{\mathrm{b}}/t$, producing only weak local distortion. At larger $U_{\mathrm{b}}/t$ the impurity becomes fully self-trapped: a repulsive impurity expels one bath particle and forms a quantized vacancy defect, while an attractive impurity binds extra one boson into a particle defect. Both MI-side defects exhibit short-range density distortion and vanishing impurity winding, in stark contrast to the extended polaronic dressing in the SF regime.